OpenStudy (anonymous):

Find the first four terms of the multiplication pattern given by the formula. 8 • 3^n – 1

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

does n starts at 0 or 1 ?

OpenStudy (anonymous):

Uhh 1?

OpenStudy (anonymous):

n=1 ---> 8*3^1-1 = 24-1=23 n=2 ---> 8*3^2-1 = 8*9-1=72-1=71 n=3 ---> 8*3^3-1 = 8*27-1=216-1=215 :)

OpenStudy (anonymous):

23, 47 ,71, 95 ?

OpenStudy (anonymous):

Wouldn't it be that..

OpenStudy (anonymous):

@texaschic101 is that right?

OpenStudy (anonymous):

you are doing: 8*3*n-1 you asked for 8*3^n-1

OpenStudy (anonymous):

right..

OpenStudy (anonymous):

So your saying the first terms are: 23,71,215 correct?

OpenStudy (anonymous):

yes, for this formula 8*3^n-1, with n starting at 1

OpenStudy (texaschic101):

correct ^^

OpenStudy (texaschic101):

if you want the first term sub in 1 for n....second term is subbing 2 for n and so on...

OpenStudy (anonymous):

And the last one is

OpenStudy (anonymous):

I need four so.

OpenStudy (texaschic101):

sub in 4 for n and solve

OpenStudy (anonymous):

haha, i had read three terms!! you can find fourth one using n=4 :)

OpenStudy (anonymous):

647?

OpenStudy (anonymous):

Is that the last term.

OpenStudy (texaschic101):

yes it is

OpenStudy (anonymous):

\[8 × 3^4 - 1 = 647\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw! :D

OpenStudy (texaschic101):

just remember that the n stands for the number of the term you want to find.

OpenStudy (anonymous):

Greg mind giving texas a medal.

OpenStudy (texaschic101):

I didn't really do anything...medal is not necessary

OpenStudy (anonymous):

sure! im new here, dont know exactly how this works...

OpenStudy (anonymous):

Still deserve it.